OpenStudy (anonymous):

NEEDED ASAP solve: -2(2x + 5) - 3 = -3(x - 1) 16 -16 -10 -13

OpenStudy (johnweldon1993):

Step 1) distribute the -2 into those parenthesis...and do the same with that -3 Step 2) combine like terms Step 3) solve for 'x'

OpenStudy (anonymous):

so negative 13?

OpenStudy (johnweldon1993):

Not quite...care to try again?

OpenStudy (johnweldon1993):

-2(2x + 5) - 3 = -3(x - 1) -4x - 10 - 3 = -3x + 3 -4x - 13 = -3x + 3 -x - 13 = 3 -x = 16 x = -16

OpenStudy (anonymous):

ooooh